OpenStudy (anonymous):

-4x - 6 = x^2 ( i think u have to use the qua. formula)

OpenStudy (radar):

It is going to be a complex number with both real and imaginary components.\[x ^{2}-4x+6=0\] using completing the square technique.\[x ^{2}-4x+4=-2\] \[(x-2)^{2}=-2\] \[x-2=\sqrt{-2}\] Taking the square root of both sides \[x=2+\sqrt{-2}\] \[x=2+\]

OpenStudy (anonymous):

? wats the answer?

OpenStudy (radar):

You can also use the i which is the\[\sqrt{-1}\] getting the final answer\[x=2+\sqrt{2}i\]

OpenStudy (anonymous):

oh i didnt learn imaginary #'s yet :P im in 8th grade algebra

OpenStudy (radar):

You may then leave it x=2+sqrt -2. The quadratic formula would of also worked.

OpenStudy (anonymous):

hmmm the answer key says 2+sqrt10 and i got 4+sqrt10 :P

OpenStudy (radar):

Let me check my work, no I have worked it using completing the square and completing the square with same solution. If I made an error maybe someone will point it out. (please) lol

OpenStudy (radar):

*i meant using the quadratic formula.

OpenStudy (radar):

elecengr was here earlier, maybe he can come up with the "approved" solution.

OpenStudy (anonymous):

how did u go from the first step to the second step?

OpenStudy (anonymous):

orig. problem to "first step"

OpenStudy (anonymous):

sry i gtg thanks for helping me (:

OpenStudy (radar):

That is the problem, I should of had a +4x, Please disregard the rest and start with\[x ^{2}+4x+6=0\]

OpenStudy (radar):

Do you see how I got that?

OpenStudy (radar):

added 4x + 6 to both sides.

OpenStudy (radar):

Now complete the square by.\[x ^{2}+4x + 4=-2\] subtracting 2 from both sides. \[(x+2)^{2}=-2\] Now take the sq root of both sides\[x+2=\sqrt{-2}\] \[x=-2+\sqrt{-2}\]

OpenStudy (radar):

Not the school book answer but different. Hope this helps.